# One Tool Is Enough: Reinforcement Learning for Repository-Level LLM Agents

### v3

@@ Appendix E Case Study
In this section, we present the full trajectory of RepoNavigator on astropy_astropy-12907 from SWE-bench_Verified in Fig. E. We apply the default tool-calling prompt template of verl (shen2024llm) and present an example. The system prompt and user prompt are presented in Fig. E. Noted, we do not present any process restrictions in our prompt, encouraging RepoNavigator to plan, call tools, and make decisions full-automatically. This is distinct with Agentless (which has a fixed workflow), LocAgent (which predefines a specific step-by-step workflow in its system prompt), CoSIL and RepoSearcher (which is half-automatic because some forced steps are added to the workflow besides the automatic multi-turns tool-calling conversations).

### v4

@@ Appendix E Case Study
In this section, we present the full trajectory of RepoNavigator on astropy_astropy-12907 from SWE-bench_Verified in Fig. E. We apply the default tool-calling prompt template of verl (Shen, 2024) and present an example. The system prompt and user prompt are presented in Fig. E. Noted, we do not present any process restrictions in our prompt, encouraging RepoNavigator to plan, call tools, and make decisions full-automatically. This is distinct with Agentless (which has a fixed workflow), LocAgent (which predefines a specific step-by-step workflow in its system prompt), CoSIL and RepoSearcher (which is half-automatic because some forced steps are added to the workflow besides the automatic multi-turns tool-calling conversations).